OpenStudy (anonymous):

URGENT HELP PLEASE!!!! Updated Batting Averages are computed using the formula: A = quantity of B plus K all over G A = batting average B = times at bat multiplied by player’s seasonal average K = hits in the new game G = times at bat plus times at bat in a new game Solve A = quantity of B plus K all over G for K. Please show your work. (6 points) Use your new formula and the values given in the table below to find the player’s hits in a new game (K). Round your answer to the nearest whole number. (4 points)

OpenStudy (anonymous):

Batter A B G K George 0.319 10.7 46 Jose 0.320 11.4 45 Dale 0.298 12.3 52 Michael 0.337 15.2 48

OpenStudy (ash2326):

George 0.319 10.7 46 Are these values A, B and G?

OpenStudy (anonymous):

yes.

OpenStudy (ash2326):

OK, it's given in the question \[A=\frac{B+K}{G}\] Can you solve this for K?

OpenStudy (ash2326):

OK, I'll guide you. First step multiply both sides by G, would you do that?

OpenStudy (anonymous):

both sides as in A and B?

OpenStudy (ash2326):

\[G\times A=G \times \frac{B+K}{G}\]

OpenStudy (anonymous):

george: 14.674 jose: 14.1 dale: 15.496 michael: 16.176

OpenStudy (ash2326):

Nope, I didn't mean the values. I meant in this formula: \[G\times A=G \times \frac{B+K}{G}\] Just simplify this

OpenStudy (anonymous):

all i need to do is solve K for all of them

OpenStudy (ash2326):

\[G\times A=\cancel G \times \frac{B+K}{\cancel G}\] \[GA=B+K \] now find the value of K from this?

OpenStudy (anonymous):

K thanks

OpenStudy (ash2326):

Did you find the value of K?

OpenStudy (anonymous):

Yeah, i guess.

OpenStudy (ash2326):

What did you find k as?

OpenStudy (anonymous):

im in the process of figuring that out

OpenStudy (ash2326):

OK, try and if you get stuck. I'll help you

OpenStudy (anonymous):

george:25.374 jose:25.5 dale:27.796 michael:31.376

OpenStudy (ash2326):

How did you find these?

OpenStudy (anonymous):

GA=B+K

OpenStudy (ash2326):

First find K in terms of GA and B

OpenStudy (anonymous):

Now I Am Very Confused!

OpenStudy (ash2326):

Suppose you have \[6=4+x\] find x from this? how'd you do this?

OpenStudy (anonymous):

x=2 6=4+2 i subtracted 4 from 6 and got 2

OpenStudy (ash2326):

So here also do the same find K \[GA=B+K\]

OpenStudy (anonymous):

now i am confused again. you have to put it in numbers not letters

OpenStudy (ash2326):

No.s and variables work the same Subtract both sides by B !

OpenStudy (anonymous):

To late time is up on my exam. but thanks for the help.